# Passage-traversing optimal path planning with sampling-based algorithms

### v1

@@ 3.1 Passage-traversing optimal path planning (PTOPP) problems
Consider an optimal path planning (OPP) problem from the start position $\mathbf{x}_{0}$ to the goal position $\mathbf{x}_{g}$. The configuration space $\mathcal{X}$ is populated with $m$ obstacles $\mathcal{O}_{1},...,\mathcal{O}_{m}$ which constitute the obstacle region $\mathcal{X}_{obs}=\cup_{i=1}^{m}\mathcal{O}_{i}$. Both $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$ are in the obstacle-free space $\mathcal{X}_{free}=\text{cl}(\mathcal{X}\backslash\mathcal{X}_{obs})$ where cl()($\cdot$)( ⋅ ) denotes the closure of a set. Our development is in the most common workspace in robotics, $\mathcal{X}\subset\mathbb{R}^{2}$ and $\mathbb{R}^{3}$. Formally, a collision-free path is a continuous function $\sigma:[0,1]\mapsto\mathcal{X}_{free}$ and the path length is defined as
@@
where $r$ is the ratio $r=$Len($\sigma_{1}$) /// Len($\sigma_{1}|\sigma_{2}$).
PTOPP problems are a class of OPP that explicitly embed the path’s passage traversal status into path cost formulation. Though a unified passage definition is unavailable, a passage $\mathcal{P}$ is a subset of $\mathcal{X}_{free}$ constructed by a pair of obstacles under some conditions. Two relations between a path $\sigma$ and $\mathcal{P}$ exist: $\sigma$ traverses $\mathcal{P}$ or otherwise. Cases where $\sigma$ partially or entirely lies in the passage ($\sigma\subset\mathcal{P}$) are trivial. Passages passed by $\sigma$ from the start $\sigma(0)$ to a path point $\sigma(\tau)$ forms an ordered list $P_{\sigma}(\tau)=\{\mathcal{P}_{i},...,\mathcal{P}_{k}\}$. $P_{\sigma}(\tau,i)$ indexes the $i$-th passage in $P_{\sigma}(\tau)$. PTOPP has the cost $f_{c}(\cdot)$ as a function of $P_{\sigma}$. For example, narrow passages passed by $\sigma$ largely determine its accessible free space. The minimum passage width in $P_{\sigma}$ thus can be employed in $f_{c}(\cdot)$ to avoid passing narrow passages. Such requirements are important when the desired path clearance cannot be prespecified properly, e.g., in deformable object manipulation and swarm navigation in obstacle-dense environments (Huang et al., 2024a; Mao et al., 2024, 2025). To sum up, PTOPP aims to find a path $\sigma^{*}$ such that $f_{c}(\sigma^{*})=\min\{f_{c}(\sigma):\sigma\text{ is feasible}\}$ where $f_{c}(\cdot)$ is a function of $P_{\sigma}$, and reports failure if no such path exists.
@@ 3.2 Passage definition with Gabriel condition
Passage detection is a fundamental prerequisite for PTOPP. Unlike RBB and its variants (Hsu et al., 2003; Sun et al., 2005) that identify narrow passages in the random sampling process, it is assumed that the obstacle distribution is known, as in most realistic scenarios. Intuitively, narrow passages are where free space significantly shrinks and becomes less accessible due to the existence of nearby obstacles. Every pair of ordered obstacles $\mathcal{O}_{i},\mathcal{O}_{j}(1\leq i<j\leq m)$ defines a passage between them. This primitive definition results in dense and unrealistic passages. The visibility condition excludes visually invalid passages blocked by other obstacles (Huang et al., 2023), but a large fraction of redundant passages remain, e.g., Figure 1. By redundant it means that a passage is not informative for evaluating accessible free space in its neighborhood. Furthermore, passage redundancy leads to higher computational loads and uncoordinated path features in related postprocessing. The goal is a sparse and informative distribution of narrow passages.
Figure 2: Passage regions between obstacles. Blue dashed segments are convex hull sides of obstacles. Passages as obstacles’ common projection shadow volume in the distance direction are in orange, which are not sensitive to relative obstacle size differences.
@@
Let $l(\mathbf{p}_{i}^{*},\mathbf{p}_{j}^{*})$ be the segment connecting $\mathbf{p}_{i}^{*}$ and $\mathbf{p}_{j}^{*}$. The recognized passage region $\mathcal{P}_{i-j}$ should cover $l(\mathbf{p}_{i}^{*},\mathbf{p}_{j}^{*})$ as the most constrained part enclosed by $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$. Starting from this observation, $\mathcal{P}_{i-j}$ grows orthogonal to $l(\mathbf{p}_{i}^{*},\mathbf{p}_{j}^{*})$ bidirectionally until it no longer simultaneously intersects with $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$. Formally, the resulting passage region equals the intersection of two obstacles’ projection shadow volumes to each other in the $l(\mathbf{p}_{i}^{*},\mathbf{p}_{j}^{*})$ direction, i.e.,
@@
where $\mathbf{e}_{i,j}$ is the unit vector of $\mathbf{p}_{j}^{*}-\mathbf{p}_{i}^{*}$ and $\mathbf{e}_{i,j}$ is its inverse, both designating the projection directions of the shadow volume SV($\cdot$). In this definition, $\mathcal{P}_{i-j}$ is insensitive to obstacle size differences as shown in Figure 2.
Not all passages given by (4) are informative in accessible free space characterization. The validity of $\mathcal{P}_{i-j}$ is examined in the presence of other obstacles. To address the passage redundancy in the visibility condition, the extended visibility condition is proposed to enforce more strict and thorough passage filtering in (Huang et al., 2024a). This criterion is inspired by investigating how robot motions are restricted inside passages. Specifically, the maximum free space limited by $l(\mathbf{p}_{i}^{*},\mathbf{p}_{j}^{*})$ is the closed disc $\mathcal{R}_{i,j}$ with $l(\mathbf{p}_{i}^{*},\mathbf{p}_{j}^{*})$ as its diameter. If any other obstacle $\mathcal{O}_{k}$ intersects with $\mathcal{R}_{i,j}$, the free space $\mathcal{R}_{i,j}$ shrinks and $\mathcal{O}_{k}$ forms passages with $\mathcal{O}_{i}$ or $\mathcal{O}_{j}$ narrower than $\mathcal{P}_{i-j}$, making $\mathcal{P}_{i-j}$ invalid. This condition is expressed as
@@
In 3D space, the most common situation where obstacles are placed on a base ground rather than floating up in midair is of interest. An interval $[\underline{h},\overline{h}]$ is used to designate the valid height range of a spatial passage. In (Huang et al., 2024a), the height interval of $\mathcal{P}_{i-j}$ is set between the base (zero height) and the lower obstacle height between $\mathcal{O}_{i},\mathcal{O}_{j}$, i.e., $\underline{h}=0,\overline{h}=\min(h_{i},h_{j})$ where $h_{i}$ and $h_{j}$ are heights of $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$, respectively. Simple as it is, this representation does not capture all narrow passages in space and only covers those on the base. As illustrated in Figure 4, a passage in 3D space exists until some lower obstacle blocks it. Therefore, the lower height bound $\underline{h}$ is the largest height among those obstacles making $\mathcal{V}(\cdot)$ false. After introducing the valid height interval, the detection condition becomes
@@
$\overline{h}=\min(h_{i},h_{j})$ is the upper height bound of $\mathcal{P}_{i-j}$. If there exists any obstacle $\mathcal{O}_{k}$ blocking $\mathcal{P}_{i-j}$ as in (5) and having a height larger than $\overline{h}$, $\mathcal{P}_{i-j}$ is invalid. Otherwise, $\mathcal{P}_{i-j}$ is valid and its lower height bound $\underline{h}$ equals the largest height of obstacles invalidating $\mathcal{V}(\mathcal{P}_{i-j})$,
@@
The condition in (5) expresses the proximity of $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$ as in the Gabriel graph (Gabriel and Sokal, 1969; Norrenbrock, 2016) that describes point nearness in the Euclidean plane. It is thus termed the Gabriel condition herein. $\mathcal{P}_{i-j}$ acts as the edge between $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$. Gabriel graph is a subgraph of the Delaunay graph, the dual graph of the Voronoi diagram (Mark et al., 2008). Although there exist other proximity graphs, e.g., the relative neighborhood graph (Toussaint, 1980; Melchert, 2013) and the nearest neighbor graph (Eppstein et al., 1997; Prokhorenkova and Shekhovtsov, 2020), as more selective subgraphs of the Gabriel graph, the Gabriel graph is more suitable for passage detection in robotics contexts because of its physical interpretation above. To the best of our knowledge, Gabriel graphs and other proximity graphs have not been applied to passage detection. Their characterization capacity of proximity relations is quite instrumental for narrow passage identification. Since only relative positions of obstacles are used, they also have better generality than explicit conditions, such as preset passage width limits in RBB.
@@ 3.3 Passage detection with Delaunay graph
where ¯$\overline{\mathcal{R}}_{i,j}=\mathcal{R}_{i,j}\cup\mathcal{P}_{i-j}$. Operations on obstacles, e.g., computation of $l(\mathbf{p}_{i}^{*},\mathbf{p}_{j}^{*})$ and $\mathcal{P}_{i-j}$, take constant time. To find all valid passages, a direct traversal check of (8) over all obstacle pairs consumes $O(m^{3})$ time. The Gabriel graph is obtainable in linear time by checking the associated Delaunay graph computable in $O(m\log m)$ time for $m$ points (Norrenbrock, 2016). However, both obstacles and passages are volumetric rather than points in reality. The graph connected by passages need not be isomorphic to the Gabriel graph with obstacles represented by simple nodes, e.g., centroids, due to obstacle sizes (see Figure 5). Existing algorithms are thereby not directly applicable.
1 Input obstacles $\mathcal{O}_{1},...,\mathcal{O}_{m}$ as polygons;
@@ 10 Append $\mathcal{P}_{i-j}$ to the passage list;
1 Sort $\mathcal{O}_{1},...,\mathcal{O}_{m}$ in descending order of height;
@@ 9 Assign $\mathcal{P}_{p-q}$’s lower height bound as $\underline{h}=h_{i}$;
                  10 Label $\mathcal{P}_{p-q}$’s height internal processed;
@@ 13 $N_{k}\leftarrow N_{i}\cup N_{j}$;
                  15 Append $\mathcal{P}_{i-j}$ to the passage list with its height interval being $[0,h_{i}]$;
@@
To efficiently reduce detection complexity, the proximity of Delaunay graph nodes is leveraged. An obstacle $\mathcal{O}_{i}$ is represented as its centroid $\mathbf{c}_{i}$. The Delaunay graph $\mathcal{DG}(C_{c})$ is computed from the centroid set $C_{c}=\{\mathbf{c}_{1},\mathbf{c}_{2},...,\mathbf{c}_{m}\}$. Although obstacle geometries vary, the geodesic distance $d_{gd}(\mathbf{c}_{i},\mathbf{c}_{j})$ between $\mathbf{c}_{i}$ and $\mathbf{c}_{j}$ in $\mathcal{DG}(C_{c})$, i.e., the number of edges in the shortest path connecting them, is small and bounded if $\mathcal{P}_{i-j}$ is valid. Hence, it suffices to examine limited neighbors of $\mathcal{O}_{i}$. The smallest geodesic distance $k_{gd}$ to be checked is
@@
$k_{gd}$ is affected by obstacle geometries and distributions. In point sets, $k_{gd}=1$ to check if an edge incident to $\mathbf{c}_{i}$ is in the Gabriel graph. As obstacles are simplified as points, $k_{gd}=1$ is no longer sufficient to find all $\mathcal{O}_{j}$ such that $\mathcal{P}_{i-j}$ is valid, e.g., Figure 5. In practice, for random obstacle distributions under limited size variations, $k_{gd}$ is small and assigned as two in our implementation (see Appendix A.1 for analysis). The expected degree of a random point in the Delaunay graph is at most six (Mark et al., 2008). Therefore, the number of edges to be examined is linear in the obstacle number. For a given pair of $\mathcal{O}_{i},\mathcal{O}_{j}$, (8) is checked only with obstacles $\mathcal{O}_{k}$ of a geodesic distance to $\mathcal{O}_{i}$ or $\mathcal{O}_{j}$ not greater than $k_{gd}$. The number of such $\mathcal{O}_{k}$ is also a constant on average. The overall passage detection algorithm using Delaunay graph of obstacles is outlined in Algorithm 1 and has $O(m\log m)$ time complexity.
Instead of directly using proximity graphs in 3D space, determining spatial passages with height intervals fits into the randomized incremental construction process of planar Delaunay graphs. Obstacles are first sorted in descending order of height. When computing the Delaunay graph, obstacles are inserted in the sorted order and passages are updated locally near the newly added obstacle $\mathcal{O}_{i}$. In particular, the update consists of two parts. First, the validity of previously detected passages that contain obstacles with geodesic distance $d_{gd}\leq k_{gd}$ to $\mathcal{O}_{i}$ is examined. Second, new passages incident to $\mathcal{O}_{i}$ and obstacles with $d_{gd}\leq k_{gd}$ to $\mathcal{O}_{i}$ are checked. This loop ends until all obstacles are added (see Algorithm 2). The core routine of passage detection is limited to the neighborhood of the newly added obstacle. Only a constant number of previously detected and new passages are processed in each iteration. Therefore, the overall detection procedure preserves the complexity of Delaunay graph construction as $O(m\log m)$ despite extra operations conducted.

### v2

@@ 3.1 Passage-traversing optimal path planning (PTOPP) problems
Consider an optimal path planning (OPP) problem from the start position $\mathbf{x}_{0}$ to the goal position $\mathbf{x}_{g}$. The configuration space $\mathcal{X}$ is populated with $m$ obstacles $\mathcal{O}_{1},...,\mathcal{O}_{m}$ which constitute the obstacle region $\mathcal{X}_{obs}=\cup_{i=1}^{m}\mathcal{O}_{i}$. Both $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$ are in the obstacle-free space $\mathcal{X}_{free}=\text{cl}(\mathcal{X}\backslash\mathcal{X}_{obs})$ where cl()($\cdot$) denotes the closure of a set. Our development is within the common workspace, $\mathcal{X}\subset\mathbb{R}^{2}$ and $\mathbb{R}^{3}$. Formally, a collision-free path is a continuous function $\sigma:[0,1]\mapsto\mathcal{X}_{free}$ and the path length is defined as
@@
where $\sigma_{1}|\sigma_{2}$ is the concatenated path of $\sigma_{1}$ and $\sigma_{2}$ satisfying $\sigma_{1}(1)=\sigma_{2}(0)$. The ratio r=r=Len($\sigma_{1}$) // Len($\sigma_{1}|\sigma_{2}$).
Figure 1: Example of dense passage distributions (orange segments) using the pure visibility condition in a laboratory for drone swarm flight, e.g., (mao2024optimal).
PTOPP problems are a class of OPP that explicitly embed the path’s passage traversal status into the path cost $f_{c}(\cdot)$. Though a unified passage definition is unavailable, a passage $\mathcal{P}$ can be defined as a subset of $\mathcal{X}_{free}$ constructed by a pair of obstacles under some conditions. A path $\sigma$ may traverse $\mathcal{P}$ or not. Cases where $\sigma$ partially or entirely lies in a passage are trivial. Passages passed by $\sigma$ from the start $\sigma(0)$ to a path point $\sigma(\tau)$ form an ordered list $P_{\sigma}(\tau)=\{\mathcal{P}_{i},...,\mathcal{P}_{k}\}$ and $P_{\sigma}(\tau,i)$ indexes the $i$-th passage in $P_{\sigma}(\tau)$. PTOPP has $f_{c}(\cdot)$ as a function of $P_{\sigma}$. For example, narrow passages passed by $\sigma$ largely determine its accessible free space. The minimum passage width in $P_{\sigma}$ can be formulated in $f_{c}(\cdot)$ to avoid passing narrow passages. Such requirements are important when a proper path clearance cannot be prespecified, e.g., in deformable object manipulation and swarm navigation in obstacle-dense environments (huang2024homotopic; mao2024optimal; mao2025tube). To sum up, PTOPP aims to find a path $\sigma^{*}$ such that $f_{c}(\sigma^{*})=\min\{f_{c}(\sigma):\sigma\text{ is feasible}\}$ with $f_{c}(\cdot)$ being a function of $P_{\sigma}$, and reports failure if no such path exists.
@@ 3.2 Passage definition with Gabriel condition
Passage detection is the prerequisite for PTOPP. Unlike RBB and its variants (hsu2003bridge; sun2005narrow) that identify narrow passages in the random sampling process, the obstacle distribution is assumed to be known. Intuitively, narrow passages are where free space significantly shrinks and becomes much less accessible due to the existence of nearby obstacles. To detect them, the visibility condition excludes visually invalid passages blocked by obstacles (huang2023deformable), whereas a large fraction of redundant passages remain, e.g., Figure 1. Redundant passages are not informative in evaluating accessible nearby free space, hence inducing uncoordinated distributions and high computational loads. Our goal is a sparse and informative distribution of narrow passages.
@@
Let $l(\mathbf{p}_{i}^{*},\mathbf{p}_{j}^{*})$ be the segment connecting $\mathbf{p}_{i}^{*}$ and $\mathbf{p}_{j}^{*}$. The passage region $\mathcal{P}_{i-j}$ should cover $l(\mathbf{p}_{i}^{*},\mathbf{p}_{j}^{*})$ as the most constrained part enclosed by $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$. Starting from this observation, $\mathcal{P}_{i-j}$ grows orthogonal to $l(\mathbf{p}_{i}^{*},\mathbf{p}_{j}^{*})$ bidirectionally until it no longer simultaneously intersects with $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$. Formally, the resulting passage region is the intersection of two obstacles’ projection shadow volumes in $l(\mathbf{p}_{i}^{*},\mathbf{p}_{j}^{*})$ direction, i.e.,
@@
where $\mathbf{e}_{i,j}$ is the unit vector of $\mathbf{p}_{j}^{*}-\mathbf{p}_{i}^{*}$ and $\mathbf{e}_{i,j}$ is its inverse, both designating the projection directions of the shadow volume SV($\cdot$). In this definition, $\mathcal{P}_{i-j}$ is insensitive to obstacle size differences as in Figure 2.
Figure 2: Passage regions between obstacles. Blue dashed segments are convex hull sides of obstacles. Passages as obstacles’ common projection shadow volume in the distance direction are in orange, which are not sensitive to obstacle size differences.
Not all passages given by (4) are informative amid multiple obstacles. The validity of $\mathcal{P}_{i-j}$ is further examined in the presence of other obstacles. To address the passage redundancy in the visibility condition, the extended visibility condition is proposed to enforce more thorough passage filtering in (huang2024homotopic). Specifically, by examining how robots are restricted inside passages, the maximum free space limited by $l(\mathbf{p}_{i}^{*},\mathbf{p}_{j}^{*})$ is the closed disc $\mathcal{R}_{i,j}$ with $l(\mathbf{p}_{i}^{*},\mathbf{p}_{j}^{*})$ as its diameter. If any other obstacle $\mathcal{O}_{k}$ intersects with $\mathcal{R}_{i,j}$, the free space shrinks and $\mathcal{O}_{k}$ forms passages with $\mathcal{O}_{i}$ or $\mathcal{O}_{j}$ narrower than $\mathcal{P}_{i-j}$, rendering $\mathcal{P}_{i-j}$ invalid. This condition is expressed as
@@
In 3D space, the most common situation where obstacles are placed on a base ground is of interest first. An interval $\underline{h},\overline{h}]$ is used to designate the valid height range of a spatial passage. In (huang2024homotopic), the height interval of $\mathcal{P}_{i-j}$ is set between the base (zero height) and the lower obstacle height between $\mathcal{O}_{i},\mathcal{O}_{j}$, i.e., $\underline{h}=0,\overline{h}=\min(h_{i},h_{j})$ where $h_{i}$ and $h_{j}$ are heights of $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$, respectively. Simple as it is, this representation does not capture all narrow passages in space and only contains those on the base. As illustrated in Figure 4, a passage in 3D space exists until some lower obstacle blocks it. Therefore, the lower height bound $\underline{h}$ is the largest height among those obstacles making $\mathcal{V}(\cdot)$ false. After introducing the valid height interval, the detection condition becomes
@@
If there exists any obstacle $\mathcal{O}_{k}$ blocking $\mathcal{P}_{i-j}$ as in (5) and of a height larger than $\overline{h}$, $\mathcal{P}_{i-j}$ is invalid. Otherwise, $\mathcal{P}_{i-j}$ is valid and its lower height bound $\underline{h}$ equals the largest height of obstacles invalidating $\mathcal{V}(\mathcal{P}_{i-j})$,
@@
The condition in (5) expresses the proximity of $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$ the same as the Gabriel graph that describes point nearness in the Euclidean space (gabriel1969new; norrenbrock2016percolation). It is thus termed the Gabriel condition here. Gabriel graph is a subgraph of the Delaunay graph, the dual graph of the Voronoi diagram (mark2008computational). Although there exist other proximity graphs, e.g., the relative neighborhood graph (toussaint1980relative; melchert2013percolation) and nearest neighbor graph (eppstein1997nearest; prokhorenkova2020graph), as more selective subgraphs, Gabriel graph is more suitable for passage detection because of its spatial interpretation. To the best of our knowledge, Gabriel graphs and other proximity graphs have not been applied to passage detection. Their characterization capacity of proximity relations is quite instrumental for narrow passage identification. Since only relative positions of obstacles are used, they have better generality than explicit criteria, such as preset passage width limits in RBB.
@@ 3.3 Passage detection with Delaunay graph
where $\overline{\mathcal{R}}_{i,j}=\mathcal{R}_{i,j}\cup\mathcal{P}_{i-j}$. Operations on obstacles, e.g., computation of $l(\mathbf{p}_{i}^{*},\mathbf{p}_{j}^{*})$ and $\mathcal{P}_{i-j}$, take constant time. To find all valid passages, a direct traversal check of (8) over all obstacle pairs consumes $O(m^{3})$ time. The Gabriel graph is obtainable in linear time by checking the associated Delaunay graph which takes $O(m\log m)$ time to compute for $m$ points (norrenbrock2016percolation). However, obstacles and passages are volumetric in reality. The obstacle graph connected by passages need not be isomorphic to the Gabriel graph with obstacles represented by simple nodes due to obstacle sizes (see Figure 5). Existing algorithms are thereby not directly applicable.
1Input obstacles $\mathcal{O}_{1},...,\mathcal{O}_{m}$ as polygons;
@@ 10 Append $\mathcal{P}_{i-j}$ to the passage list;
1Sort $\mathcal{O}_{1},...,\mathcal{O}_{m}$ in descending order of height;
@@ 9 Assign $\mathcal{P}_{p-q}$’s lower height bound as $\underline{h}=h_{i}$;
       10 Label $\mathcal{P}_{p-q}$’s height interval processed;
@@ 13 $N_{k}\leftarrow N_{i}\cup N_{j}$;
       15 Append $\mathcal{P}_{i-j}$ to the passage list with its height interval being $0,h_{i}]$;
@@
To reduce detection complexity, the proximity of Delaunay graph nodes is leveraged. An obstacle $\mathcal{O}_{i}$ is represented as its centroid $\mathbf{c}_{i}$. The Delaunay graph $\mathcal{DG}(C_{c})$ is computed from the centroid set $C_{c}=\{\mathbf{c}_{1},\mathbf{c}_{2},...,\mathbf{c}_{m}\}$. Although obstacle geometries are indeterminate, the geodesic distance $d_{gd}(\mathbf{c}_{i},\mathbf{c}_{j})$ between $\mathbf{c}_{i}$ and $\mathbf{c}_{j}$ in $\mathcal{DG}(C_{c})$, i.e., the number of edges in the shortest path connecting them, is small and bounded if $\mathcal{P}_{i-j}$ is valid. Hence, it suffices to examine limited neighbors of $\mathcal{O}_{i}$. The smallest geodesic distance $k_{gd}$ to be checked is
@@
$k_{gd}$ is affected by obstacle geometries and distributions. In point sets, $k_{gd}=1$ to check if an edge incident to $\mathbf{c}_{i}$ is in the Gabriel graph. As obstacles are simplified as points, $k_{gd}=1$ is no longer sufficient to find all $\mathcal{O}_{j}$ such that $\mathcal{P}_{i-j}$ is valid, e.g., Figure 5. In practice, for random obstacle distributions under limited size variations, $k_{gd}$ is small and assigned as two in our implementation (see Appendix A.1 for analysis). For a given pair of $\mathcal{O}_{i},\mathcal{O}_{j}$, (8) is checked only with obstacles $\mathcal{O}_{k}$ of $d_{gd}$ to $\mathcal{O}_{i}$ or $\mathcal{O}_{j}$ not greater than $k_{gd}$. Let such a neighbor list of $\mathcal{O}_{i}$ be $N_{i}(k_{gd})=\{\mathcal{O}_{k}|d_{gd}(\mathcal{O}_{i},\mathcal{O}_{k})\leq k_{gd}\}$. Then passages incident to $\mathcal{O}_{i}$ are detected between $\mathcal{O}_{j}\in N_{i}(k_{gd})$ and
$\begin{split}\mathcal{V}(\mathcal{P}_{i-j})&=\text{False}\;\;\text{if}\;\;\mathcal{O}_{k}\cap\overline{\mathcal{R}}_{i,j}\neq\emptyset,\\ \exists\,\mathcal{O}_{k}&\in N_{i}(k_{gd})\cup N_{j}(k_{gd}).\end{split}$ (10)
The expected degree of a Delaunay graph node is at most six (mark2008computational). Therefore, the number of edges to be examined is linear in the number of obstacles and the cardinality $|N_{i}(k_{gd})|$ is a constant on average. The overall $O(m\log m)$ passage detection algorithm using Delaunay graph of obstacles is outlined in Algorithm 1.
Instead of using 3D proximity graphs, determining spatial passages with height intervals fits into the randomized incremental construction process of planar Delaunay graphs. Obstacles are first sorted in descending order of height. When computing the Delaunay graph, obstacles are inserted in the sorted order and passages are updated locally near the newly added obstacle $\mathcal{O}_{i}$. The update consists of two parts. First, the validity of previously detected passages that contain obstacles in $N_{i}(k_{gd})$ is examined. Second, new passages between $\mathcal{O}_{i}$ and obstacles in $N_{i}(k_{gd})$ are checked. This loop ends until all obstacles are added (see Algorithm 2). As such, the core routine of passage detection is limited to the neighborhood of the newly added obstacle. Only a constant number of previously detected and new passages are processed in each iteration. Therefore, the overall detection procedure preserves the complexity of Delaunay graph construction as $O(m\log m)$.